OpenStudy (anonymous):

sin^2Θ - cos^2Θ = sin Θ SIMPLIFY. pls help!!!

OpenStudy (cwrw238):

cos^2Θ = 1 - sin^2Θ so sin^2Θ - (1 - sin^2Θ) = sinΘ 2sin^2Θ - 1 = sinΘ do you want to solve this for sinΘ ?

OpenStudy (anonymous):

u just need to simply it, thats all, u dont have to technically "solve it"

OpenStudy (anonymous):

so what would u do from there

OpenStudy (cwrw238):

well subtract sin Θ from both sides; 2 sin^2Θ - sinΘ - 1 = 0 factorise: (2sinΘ + 1)(sinΘ - 1) = 0 and thats as simple as you'll get it